Procedure: This experiment will utilize the MicroLab 522 interface, a pH probe and a drop counter to perform the studies. 1. Preform calculation to determine the mass of solid NaOHl which must be weighed to prepare 250mL of a 0.2M NaOH solution. Preform calculation to determine the mass of solid Benzoic Acid which must be weighed to prepare 100mL of a 0.1M Benzoic Acid solution. Calibrate your pH Probe using a 3 point calibration with pH 4, 7 and 10 buffers 2. 3. Set the drop counter at an optimum drop rate and use the conversion factor of 0.034 mL per drop to convert to volume titrant used. 4.
Answer #1

1

250 mL 0.2 M NaOH = 250 x 0.2 = 50 mmol NaOH

Molar mass of NaOH = 1 x 23 + 1 x 16 +1 x 1 = 40 g/mol

Thus the mass of 1 mol NaOH = 40 g

Therefore the mass of 50 mmol or 50 x 10-3 mol NaOH = 40 x 50 x 10-3 = 2 g

Therefore the mass of solid NaOH must be weighed for making 250 mL 0.2 M NaOH = 2 g

---------------------------------------------------------

2

100 mL 0.1 M benzoic acid = 100 x 0.1 = 10 mmol benzoic acid

Molecular formula of benzoic acid = C6H5COOH = C7H6O2

Therefore the molar mass of benzoic acid = 7 x 12 + 6 x 1 + 2 x 16 = 84 + 6 + 32 = 122 g/mol

Thus the mass of 1 mol benzoic acid = 122 g

Therefore the mass of 10 mmol or 10 x 10-3 mol benzoic acid = 122 x 10 x 10-3 g = 1.22 g

Therefore the mass of solid benzoic acid must be weighed for making 100 mL 0.1 M benzoic acid = 1.22 g